# SEDM: Scalable Self-Evolving Distributed Memory for Agents

### v1

@@ 4.1 Experimental Setup
Experiments are conducted on two widely used benchmark datasets: FEVER(Thorne et al., 2018) and HotpotQA(Yang et al., 2018b).HotpotQA is a large-scale question-answering benchmark designed to evaluate the ability of systems to perform multi-hop reasoning across diverse natural language inputs. FEVER is a fact-checking dataset that provides human-written claims about Wikipedia entities, each labeled as Supported, Refuted, or NotEnoughInfo. Both datasets present significant challenges for testing long-term reasoning and memory utilization in language agents.
Baselines.
@@
The proposed SEDM is compared with the following baselines: (1) No Memory: the model only relies on the query input without any memory augmentation, serving as the basic performance reference; (2) G-Memory(Zhang et al., 2025): a memory-augmented method that stores all past information in a global memory pool and retrieves by similarity search. Although effective, it incurs high inference cost due to the large number of prompt tokens; (3) SEDM (ours): our scalable self-evolving distributed memory, which introduces memory scheduling and selection mechanisms to maintain a compact and adaptive working set, thereby balancing performance and efficiency.
Configurations.
@@
All experiments run on the same backbone LLM (GPT-4o-mini) OpenAI (2023). Dense retrieval is handled by ALL-MINILM-L6-V2 Reimers and Gurevych (2019), which embeds both knowledge snippets and queries for similarity search. On the evaluation side, we use FEVER accuracy for fact-checking and HotpotQA exact-match (EM) for multi-hop QA. Efficiency is tracked by counting prompt and completion tokens consumed during inference. To ensure fair comparisons, every method is granted the same memory budget; the proposed SEDM does not expand this budget, but instead adaptively schedules which entries are kept in memory.
4.2 Result Analysis
@@ 4.4 Cross-Domain Evaluation
To further assess the generalization ability of SEDM across domains, we conduct a cross-domain experiment in which memory is collected on one dataset and evaluated on another. Table 3 reports the results on both FEVER and HotpotQA.
@@
Collect $\downarrow$ / Test $\rightarrow$ Fever HotpotQA
Fever 66 41
HotpotQA 64 39
The results highlight two key findings. First, in-domain settings (diagonal entries) achieve the strongest performance, with scores of 66 on FEVER and 39 on HotpotQA, confirming that SEDM effectively distills and leverages domain-specific memory. Second, cross-domain transfer (off-diagonal entries) reveals an interesting asymmetry. When memory distilled from FEVER is applied to HotpotQA, the model achieves a score of 41, which even surpasses the in-domain HotpotQA result (39). This indicates that factual knowledge captured in FEVER has strong generalization capacity and can support reasoning tasks that require grounding in factual claims. By contrast, memory distilled from HotpotQA transfers less effectively to FEVER (64 vs. 66), suggesting that multi-hop reasoning knowledge is less directly reusable for fact verification. Overall, these findings demonstrate that SEDM not only excels in in-domain scenarios but also exhibits promising cross-domain generalization, particularly when transferring from simpler fact verification tasks to more complex multi-hop reasoning tasks.
@@ 5 Conclusion
This paper introduces SEDM, Scalable Self-Evolving Distributed Memory, which transforms memory in multi-agent systems from a passive repository into an adaptive and verifiable component by integrating SCEC-based admission, self-scheduling refinement, and cross-domain knowledge diffusion. Through this principled design, SEDM addresses the challenges of noise accumulation, uncontrolled growth, and weak generalization that limit existing methods. Experiments on FEVER and HotpotQA confirm that SEDM improves reasoning accuracy while reducing computational and token overhead, demonstrating its potential as a scalable and sustainable memory mechanism for long-term multi-agent collaboration.

### v3

@@ 4.3 Efficient analysis on Fever and HotpotQA
We further conduct Experiments on two widely used benchmark datasets: FEVER(thorne-etal-2018-fever, ) and HotpotQA(yang2018hotpotqadatasetdiverseexplainable, ).HotpotQA is a large-scale question-answering benchmark designed to evaluate the ability of systems to perform multi-hop reasoning across diverse natural language inputs. FEVER is a fact-checking dataset that provides human-written claims about Wikipedia entities, each labeled as Supported, Refuted, or NotEnoughInfo. Both datasets present significant challenges for testing long-term reasoning and memory utilization in language agents.
@@
The proposed SEDM is compared with the following baselines: (1) No Memory: the model only relies on the query input without any memory augmentation, serving as the basic performance reference; (2) G-Memory(zhang2025gmemorytracinghierarchicalmemory, ): a memory-augmented method that stores all past information in a global memory pool and retrieves by similarity search. Although effective, it incurs high inference cost due to the large number of prompt tokens; (3) SEDM (ours): our scalable self-evolving distributed memory, which introduces memory scheduling and selection mechanisms to maintain a compact and adaptive working set, thereby balancing performance and efficiency.
@@
All experiments run on the same backbone LLM (GPT-4o-mini) openai2024gpt4 . Dense retrieval is handled by ALL-MINILM-L6-V2 reimers2019sentencebert , which embeds both knowledge snippets and queries for similarity search. On the evaluation side, we use FEVER accuracy for fact-checking and HotpotQA exact-match (EM) for multi-hop QA. Efficiency is tracked by counting prompt and completion tokens consumed during inference. To ensure fair comparisons, every method is granted the same memory budget; the proposed SEDM does not expand this budget, but instead adaptively schedules which entries are kept in memory.
@@ 4.5 Cross-Domain Evaluation
To further assess the generalization ability of SEDM across domains, we conduct a cross-domain experiment in which memory is collected on one dataset and evaluated on another. Table 4 reports the results on FEVER, HotpotQA, and LoCoMo.
@@
Collect $\downarrow$ / Test $\rightarrow$ FEVER HotpotQA LoCoMo
FEVER 66 41 38.1
HotpotQA 64 39 38.6
LoCoMo 65 34 37.6
FEVER as target: The best score is achieved by FEVER itself (66); memory from HotpotQA and LoCoMo yields slightly lower scores (64 and 65), indicating that fact-verification benefits most from in-domain memory.
HotpotQA as target: The highest score is obtained by FEVER → HotpotQA (41), 2 points above the in-domain result (39). Memory from LoCoMo drops to 34, the lowest cross-domain score, suggesting that dialogue-grounded knowledge is least useful for multi-hop reasoning.
LoCoMo as target: All three sources perform within 1 point of each other (37.6–38.6). Thus, no single source dominates, and dialogue-grounded evaluation is remarkably robust to the origin of memory.
Overall, SEDM exhibits task-dependent transfer: factual-verification memory transfers surprisingly well to HotpotQA, whereas dialogue memory transfers poorly to the other two tasks. In-domain memory is not universally optimal, and the benefit of domain alignment varies significantly by task pair.
@@ 5 Conclusion
This paper introduces SEDM, Scalable Self-Evolving Distributed Memory, which transforms memory in multi-agent systems from a passive repository into an adaptive and verifiable component by integrating SCEC-based admission, self-scheduling refinement, and cross-domain knowledge diffusion. Through this principled design, SEDM addresses the challenges of noise accumulation, uncontrolled growth, and weak generalization that limit existing methods. Experiments on LoCoMo, FEVER, and HotpotQA confirm that SEDM improves reasoning accuracy while reducing computational and token overhead, demonstrating its potential as a scalable and sustainable memory mechanism for long-term multi-agent collaboration.